what is 3x+-2y=4 in slope intercept form
Found 2 solutions by josgarithmetic, Theo:
Answer by josgarithmetic(39838)

y=mx+b

Solve for y and put into general form. That is slope-intercept form.

Answer by Theo(13342)

slope intercept form is y = mx + b
m is the slope and b is the y-intercept.

start with 3x - 2y = 4
subtract 3x from both sides of the equation to get -2y = -3x + 4
divide both sides of the equation by -2 to get y = (-3/-2)x + 4/-2.
simplify to get y = (3/2)x - 2.
the slope is 3/2.
the y-intercept is -2.

using both formulas, test a couple of points to make sure they give you the same answer.

when x = 0, 3x - 2y = 4 becomes -2y = 4 which becomes y = -2.
the x = 0 y = (3/2)x - 2 becomes y = -2.

when x = 2, 3x - 2y = 4 becomes 6 - 2y = 4 which becomes -2y = -2 which becomes y = 1.
when x = 2, y = (3/2)x - 2 becomes (3/2)*2 - 2 which becomes 3 - 2 which becomes y = 1.

the two equations give you the same result.
3x - 2y = 4 is the standard form.
y = (3/2)x - 2 is the slope intercept form.